OpenStudy (anonymous):

given the equations 2x+4/3y=1 and y-9/13x=9, by what factor would you multiply the first equation so that combining the two equations would eliminate x?

OpenStudy (texaschic101):

2x + 4/3y = 1 -9/13x + y = 9 are you allowed to change the equation to where there is no decimals first ?

OpenStudy (texaschic101):

oops..I mean no fractions

OpenStudy (anonymous):

No sadly it cannot :/

OpenStudy (anonymous):

Yes, fractions can be.

OpenStudy (anonymous):

−9/26 9/26 1/2 −9/13

OpenStudy (texaschic101):

Its gonna be 9/26

OpenStudy (anonymous):

Thank you!

OpenStudy (texaschic101):

sure thing :)